OpenStudy (anonymous):

The point P with the coordinates (3,m) lies on the curve y=x^2 + kx. At P the gradient of the curve is 8. Find the values of the constants k and m.

OpenStudy (unanimoose):

The answer lies within your heart.

OpenStudy (anonymous):

Thanks?

OpenStudy (anonymous):

The point P with the coordinates (3,m) lies on the curve y=x^2 + kx. At P the gradient of the curve is 8. Find the values of the constants k and m. Knowing that \((3,m)\) is a point on the curve is the same as knowing that when \(x=3\), you have \(y=m\). So you have one equation: \[m=3^2+3k=9+3k\] The gradient of the curve at this point (when \(x=3\)) is the value of the derivative at this point, which is given to be 8. So, you also know that \[\frac{dy}{dx}=2x+k~~\implies~~8=6+k\] You can directly solve for \(k\) above, then use it to find the value of \(m\).